OpenStudy (anonymous):

Is this right? Problem: 5/2y-7 = -1 my answer: 1

OpenStudy (lgbasallote):

is this 5/(2y-7) = -1? then cross multiply 5 = -(2y-7) 5 = -2y + 7 -2y = -2 y = 1

OpenStudy (anonymous):

So, im right? :D

sam (.sam.):

yes